each of the two congruent sides of an isosceles triangle is 7cm shorter than twice the base. find the length of the three sides if perimeter of triangle is 106cm
Answer by cleomenius(959)

Let x = the base
2(2x - 7) + x = 106
4x -14 + x = 106
5x - 14 = 106
5x = 120
x = 24 cm the base.
48 - 7
41 cm One side of the triangle
41 cm + 41 cm + 24 cm = 106 cm
This does check.
Cleomenius.